OpenStudy (anonymous):

@fathomboy What is the prime factorization of 120? A. 22 • 3 • 5 B. 23 • 3 • 5 C. 24 • 3 D. 23 • 5

OpenStudy (anonymous):

are the double digits like 2^3 or are they just 23?

OpenStudy (anonymous):

because if it was in exponential form it would be, B

OpenStudy (anonymous):

like 2^3

OpenStudy (anonymous):

no one of option is factor

OpenStudy (anonymous):

the factors of 120 are 2,2,2,3,5 2^3 * 3 * 5 = 120

OpenStudy (amistre64):

how do we find prime factors? as opposed to just saying what they are for this particular case ...

OpenStudy (cwrw238):

if the number is even divide by 2 and keep doing so until you get an odd result then try 3 , 5 and subsequent prime numbers

OpenStudy (cwrw238):

fathomboy has the correct result

OpenStudy (anonymous):

keep dividing for example: 120/2=60/2=30/2=15/3=5/5=1